A 4m rod AB slides down the plane with vA = 4 m/s to the left and aA = 5 m/s2 to the right. Angle θ at this moment is 30o. Determine the angular velocity and angular acceleration at this instant of time.

Grade:11

1 Answers

ashish kumar
14 years ago

Angular velocity of AB =( VA-VB)/LAB

                                      = (4 cos60 - 0)/4

                                     = 2/4

                                    = 0.5

Insatantaneous centre of rotation is point B since it has zero velocity.

Let the angular acceleration be

Therefore   aA = (AB)

                 = 5 cos60/4

 

                    = 5/8